### v3

# Graph Transformers without Positional Encodings

Abstract
Recently, Transformers for graph representation learning have become increasingly popular, achieving state-of-the-art performance on a wide-variety of graph datasets, either alone or in combination with message-passing graph neural networks (MP-GNNs). Infusing graph inductive-biases in the innately structure-agnostic transformer architecture in the form of structural or positional encodings (PEs) is key to achieving these impressive results. However, designing such encodings is tricky and disparate attempts have been made to engineer such encodings including Laplacian eigenvectors, relative random-walk probabilities (RRWP), spatial encodings, centrality encodings, edge encodings etc. In this work, we argue that such encodings may not be required at all, provided the attention mechanism itself incorporates information about the graph structure. We introduce Eigenformer, a Graph Transformer employing a novel spectrum-aware attention mechanism cognizant of the Laplacian spectrum of the graph, and empirically show that it achieves performance competetive with SOTA Graph Transformers on a number of standard GNN benchmarks. Additionally, we theoretically prove that Eigenformer can express various graph structural connectivity matrices, which is particularly essential when learning over smaller graphs.
1 Introduction
Learning useful representations from graph data is important in a variety of domains due to the ubiquitous presence of graphs around us, ranging from molecules to social and transportation networks. It may well be argued that "graphs are the main modality of data we receive from nature" [Veličković, 2023]. In general, machine learning methods designed to learn patterns from graph data fall into two paradigms: message-passing graph neural networks (MP-GNNs), where vector messages are exchanged between neighboring nodes and updated using neural networks [Bruna et al., 2014, Gilmer et al., 2017], and more recently Transformers [Vaswani et al., 2017] –originally designed for Natural Language Processing (NLP) tasks– for graph data, which feature long-range connections between nodes for faster information exchange albeit with a loss of graph inductive biases [Dwivedi and Bresson, 2020, Kreuzer et al., 2021, Chen et al., 2022a, Ying et al., 2021, Rampášek et al., 2022, Hussain et al., 2022, Zhang et al., 2023].
In many ways, MP-GNNs and Graph Transformers have complementary strengths and weaknesses. While MP-GNNs have strong graph inductive biases due to sharing of messages within local neighborhoods, they suffer from issues such as over-squashing [Alon and Yahav, 2021, Topping et al., 2022], over-smoothing [Li et al., 2018, Oono and Suzuki, 2020] and expressivity concerns (upper bounded by the 1-Weisfeiler-Lehman isomorphism test) [Xu et al., 2019, Loukas, 2020, Morris et al., 2019]. On the other hand, Graph Transformers do not suffer from these issues, at the expense of losing graph connectivity information (forgoing locality in favor of all-to-all connectivity between nodes), possibility of over-fitting and quadratic computational and memory complexity. Due to their complementary benefits, researchers have (successfully) tried to merge the two paradigms together into one architecture with GraphGPS in Rampášek et al. [2022].
A common weakness in both methods, however, is the lack of node positional information, which has been shown to be crucial both to improve expressivity in MP-GNNs (as compared to the 1-WL test in Xu et al. [2019]), and to allow Transformers to be used with datasets with smaller graphs. Nodes in a graph lack a canonical ordering, and superficially imposing an ordering leads to learning difficulties due to the exponential growth in such possible orderings. To tackle this, positional encodings (PEs) based on Laplacian eigenvectors were proposed in Dwivedi and Bresson [2020]. Due to sign ambiguities of eigenvectors and eigenvalue multiplicities, sign and basis invariant/equivariant methods such as SignNet and BasisNet in Lim et al. [2023] were proposed. In other works [Dwivedi et al., 2022a, Ma et al., 2023], $k$-dimensional (learnable) random-walk probabilities from a node to itself are used instead of Laplacian eigenvectors to avoid the aforementioned invariance issues.
Main contributions: In light of the above background, we ask the question whether such encodings are indeed necessary at all. If there is no canonical node-ordering inherent to a graph, perhaps no such ordering is required to effectively learn patterns from graph data. Thus, in this work, we seek to circumvent the issues related to designing node/edge PEs by modifying the way node feature information is exchanged in a Graph Transformer. Specifically:
  • Instead of adding positional information to node features and letting the attention mechanism figure out the strength of connection between a pair of nodes, we factorize the attention matrix in terms of fixed node-pair "potentials" and learned frequency "importances". We posit such a factorization encodes important graph inductive biases from the frequency domain into the attention mechanism while being flexible enough to yield impressive performance across domains and objectives.
  • We utilize valuable information from both the eigenvectors and the eigenvalues as advocated in Kreuzer et al. [2021], forgoing the assumption that higher frequencies are less important in determining node-pair interactions.
  • The novel attention mechanism with built-in inductive biases allows us to train wider/deeper models for a given parameter budget due to simpler Transformer layers. Numerical experiments demonstrate the strength of our method, achieving near-SOTA performances on many common GNN benchmarks.
2 Theoretical Motivations
We first present the motivation behind the architecture of Eigenformer by briefly reviewing the semantics of the graph Laplacian, its spectrum and spectral convolution.
2.1 Graph Laplacian and its spectrum
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the unnormalized graph Laplacian $L$ is defined as
$L=D-A$ (1)
where $A$ is the adjacency matrix and $D$ is the degree matrix. By definition 1, the following is true for a unit-norm eigenvector $u_{i}$ and its corresponding eigenvalue $\lambda_{i}$:
$\lambda_{i}=u_{i}^{T}Lu_{i}=\frac{1}{2}\sum_{u,v\in\mathcal{V}}A[u,v](u_{i}[u]-u_{i}[v])^{2}=\sum_{(u,v)\in\mathcal{E}}(u_{i}[u]-u_{i}[v])^{2}$ (2)
If the eigenvectors are interpreted as "signals" over the graph, taking real scalar values at nodes, then the eigenvalues act as estimates of the "smoothness" of such signals, measuring how much the signal values differ between endpoints of edges. The smallest eigenvalue is always 0, assigning the same scalar value $u_{i}[\cdot]$ to all nodes in the same connected component of the graph.
Thus, eigenvectors of the Laplacian encode important information about node similarities (nodes resonant under a frequency tend to have similar values w.r.t its eigenvector). As such, the spectrum –in its interpretation as frequencies of resonance of the graph– discriminates between different graph structures and substructures. Figure 1 shows substructures revealed by eigenvectors for an example molecular graph from the PCQM4Mv2 dataset [Hu et al., ] containing 15 nodes. Color-codes indicate the value $u_{i}[\cdot]$ for different frequencies $\lambda_{i}$ where $i\in\{0,...,15\}$. Lower frequencies highlight larger substructures while higher frequencies emphasize more local interactions.
Figure 1: Example molecule from PCQM4Mv2 dataset: Substructures are revealed by eigenvectors
2.2 Spectral convolution
Since the (sinusoidal) eigenfunctions $f(t)=e^{2\pi ist}$, of the general Laplace operator $\Delta$, correspond to the modes of the frequency domain of the Fourier transform, the eigenvectors of the graph Laplacian $L$ can be equated to sine functions over graphs. This connection allows us to generalize the Fourier transform to arbitrary graphs.
For a signal $f\in\mathbb{R}^{|\mathcal{V}|}$ on an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with Laplacian $L=U\Lambda U^{T}$, where $U$ is the orthonormal matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues, we can compute its Fourier transform as $s=U^{T}f$, and its inverse as $f=Us$. Thus, its graph convolution $\star_{\mathcal{G}}$, with a filter $h$ can be computed as:
$f\star_{\mathcal{G}}h=U(U^{T}f\odot U^{T}h)=(U\text{diag}(w_{h})U^{T})f$ (3)
where the last equality is true when the filter $h$ can be learnt in a non-parametric way by representing and optimizing directly in the spectral domain: $w_{h}=U^{T}h$. Such a filter, however, may not satisfy sound graph convolution properties like locality and translational equivariance, and it is useful to parameterize $w_{h}$ in terms of eigenvalues $\Lambda$ of the Laplacian, for example, as a polynomial $p(\Lambda)$, which ensures commutativity with the Laplacian, restoring the required properties:
$f\star_{\mathcal{G}}h=(Up(\Lambda)U^{T})f=p(L)f$ (4)
Thus, a spectral convolution on the graph can be defined in terms of a polynomial of the Laplacian $L$. The degree $k$, of the polynomial determines the $k$-hop neighborhood over which the convolution is performed.
3 Model Architecture
In this section, we introduce our novel spectrum-aware attention (SAA) mechanism that builds on the spectral convolution theory discussed in 2.2 and demonstrate its use in the proposed Eigenformer architecture.
Note: We use the normalized variant of the graph Laplacian, $L_{norm}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ that admits similar properties and interpretations of the spectrum as that of $L$, but has better numerical stability due to a bounded spectrum: all eigenvalues of $L_{norm}$ lie between 00 and $2$. Moreover, the normalized Laplacian counters the effect of "heavy" nodes (nodes with high degrees) in influencing information propagation through the graph. Also, for directed graphs, we first make the edges undirected before calculating $L_{norm}$ so that its spectrum remains real.
3.1 Attention using the Laplacian spectrum
For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ we define the attention weight between two nodes $i$ and $j$ in $\mathcal{G}$, through the following equations:
$\sigma_{k}[i,j]=u_{k}[i]\cdot u_{k}[j]$ (5)
$\alpha[i,j]=softmax_{j\in\mathcal{V}}\Bigl{[}\phi_{1}\Bigl{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\phi_{2}(\lambda_{k})\Bigl{)}\Bigl{]}$ (6)
where $u_{k}\in\mathbb{R}^{|\mathcal{V}|}$ is the eigenvector corresponding to the eigenvalue $\lambda_{k}\in\mathbb{R}$ of $L_{norm}$, $\sigma_{k}\in\mathbb{R}$ can loosely be interpreted as the "potential" between the nodes under the frequency $\lambda_{k}$ (see Green’s function from electrostatic theory, Chung and Yau [2000]), $\phi_{1}:\mathbb{R}\rightarrow\mathbb{R}$ and $\phi_{2}:\mathbb{R}\rightarrow\mathbb{R}$ are arbitrary real functions and $\alpha\in\mathbb{R}$ is the final attention weight. The connection to spectral convolution can be inferred by noting that $\alpha=softmax[\phi_{1}(U\phi_{2}(\Lambda)U^{T})]$ where $U=[u_{1},...,u_{|\mathcal{V}|}]$.
Propositions 1 and 2 below justify the choice of $\sigma$ and $\alpha$ rigorously.
Proposition 1
For any $n\in\mathbb{N}$, consider the adjacency matrix $A$ drawn from the set of adjacency matrices of n-node undirected graphs, $\mathbb{G}_{n}\subset\{0,1\}^{n\times n}$. Further, let $L_{norm}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{norm}$ be the normalized graph Laplacian of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$, adjacency matrix $A$, eigenvalues $\lambda_{k}$ and eigenvectors $u_{k}$, for $k\in\{1,...,n\}$. Then, we have the following approximations up to an arbitrary small error, $\epsilon$:
  1. 1.
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx\sum_{k=0}^{m}\theta_{k}A_{norm}^{k}[i,j]$ for $m\in\mathbb{Z}$
  2. 2.
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx f(SPD[i,j])$
for suitable functions $\phi_{1}$ and $\phi_{2}$, where $SPD[i,j]$ is the shortest path distance between nodes $i$ and $j$ and $f$ is any continuous function.
Proof: The proof of (1.) utilizes the diagonalizability of the matrices $A_{norm}$ and $L_{norm}$, allowing us to derive the analytic form of $\phi_{1}$ and $\phi_{2}$. For (2.), we show that $\phi_{2}$ can be chosen to yield a valid (possibly many-to-one) function that maps $\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})$ to $f(SPD[i,j])$ for all node-pairs $(i,j)$ in a graph $\mathcal{G}$ in the family of graphs indexed by $\mathbb{G}_{n}$, and we define $\phi_{1}$ to be that piecewise-linear map. The detailed analysis can be found in Appendix A.1).
Proposition 1 states that the attention mechanism defined in 6 can capture (i) shortest path distances between all pairs of nodes and (ii) all weighted combinations of powers of the normalized adjacency matrix $A_{norm}$ allowing us to perform aggregations over any $k-$hop neighborhood where $k\in\mathbb{Z}$. The function $\phi_{2}$ quantifies the "importance" of the frequency $\lambda_{k}$ in the sum in 6. Traditional distance metrics like the biharmonic distance [Lipman et al., 2010] or the diffusion distance [Coifman and Lafon, 2006, Bronstein et al., 2017], weigh smaller frequencies more heavily, assuming that higher frequencies play a smaller role in the measurement of relative distance between nodes. We make no such assumption, rather allowing the dependence on the frequency to be learnt from data by parameterizing the function $\phi_{2}$ appropriately.
We now show that the proposed attention mechanism does not suffer from eigenvector sign/basis-invariance issues, making prior interventions like SignNet/BasisNet [Lim et al., 2023] redundant.
Proposition 2
The attention matrix $\alpha$ computed using equations 5 and 6, is invariant to the choice of the sign of the eigenvectors $u_{k}$, and in the general case, to the choice of basis of the eigenspace corresponding to the eigenvalue $\lambda_{k}$ having geometric multiplicity > 1.
Proof: The sign-invariance can be easily inferred by noting that eigenvectors $u_{k}$ appear in the computation of $\alpha$ as multiplied pairs of terms (5). The invariance to basis is due to the fact that in 6, the eigenvectors $U_{k}=[u_{k_{1}},...,u_{k_{d_{i}}}]\in\mathbb{R}^{n\times d_{i}}$ corresponding to the same eigenvalue $\lambda_{k}$ occur together as terms of the form $h(U_{k}U_{k}^{T})$, for some continuous function $h$. Since any other orthonormal basis is of the form $V_{k}=U_{k}Q$ for some orthogonal $Q\in O(d_{i})\subseteq\mathbb{R}^{d_{i}\times d_{i}}$, we have $h(V_{k}V_{k}^{T})=h(U_{k}U_{k}^{T})$, implying that the choice of basis does not affect the computation in 6.
(a) SAA: Spectrum-aware attention
(b) A layer of Eigenformer
Figure 2: Eigenformer Architecture
3.2 Eigenformer architecture
For a layer $l$ of Eigenformer, input node features $x_{i}^{l}\in\mathbb{R}^{d_{\mathcal{V}}}$ for $i=1,...,|\mathcal{V}|$ and input edge features $e_{ij}\in\mathbb{R}^{d_{\mathcal{E}}}$ for $i,j=1,...,|\mathcal{V}|$, the node feature propagation is performed as follows:
${\alpha^{lh}_{spec}}=\phi_{1}^{lh}\Bigl{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\phi_{2}^{lh}(\lambda_{k})\Bigl{)}$ (7)
${\alpha^{lh}_{feat}}=\psi\Bigl{(}\frac{(W^{lh}_{Q}x_{i}^{l})^{T}(W^{lh}_{K}x_{j}^{l})}{\sqrt{d/H}}\Bigl{)}+W^{lh}_{R}\rho\bigl{(}W^{lh}_{A}e_{ij}\bigl{)}$ (8)
$\alpha^{lh}[i,j]=softmax_{j\in\mathcal{V}}({\alpha^{lh}_{spec}}+{\alpha^{lh}_{feat}})$ (9)
[…]
$\tilde{x}_{i}^{l+1}=\hat{x}_{i}^{l+1}\odot\theta_{1}+log(1+d_{i})\cdot\hat{x}_{i}^{l+1}\odot\theta_{2}$ (11)
$x_{i}^{\prime l+1}=BN(x_{i}^{l}+\tilde{x}_{i}^{l+1})$ (12)
$x_{i}^{l+1}=BN(x_{i}^{\prime l+1}+FFN(x_{i}^{\prime l+1}))$ (13)
[…]
$\alpha_{feat}$ is an optional attention component that we introduce to study the gains that node and edge feature attention brings over solitary spectral attention $\alpha_{spec}$. We do not include edge-feature propagation due to the added complexity and marginal benefits in performance. Notably however, we found degree-scaling followed by batch-normalization to be highly effective in incorporating node-degree information. The final architecture of a layer of Eigenformer is illustrated in Figure 2(b) along with the spectrum-aware attention mechanism that it uses in Figure 2(a).
Table 1: Test performance on five benchmarks from Dwivedi et al. [2020]. Due to limited training resources, we show results for only a single training run while other baselines report mean $\pm$ standard deviation of 4 runs with different random seeds. Color codes denote top first, second, third results.
Model ZINC MNIST CIFAR10 PATTERN CLUSTER
[…]
GCN 0.367$\pm$.011 90.705$\pm$.218 55.710$\pm$.381 71.892$\pm$.334 68.498$\pm$.976
GIN 0.526$\pm$.051 96.485$\pm$.252 55.255$\pm$1.527 85.387$\pm$.136 64.716$\pm$1.553
GAT 0.384$\pm$.007 95.535$\pm$.205 64.223$\pm$.455 78.271$\pm$.186 70.587$\pm$.447
GatedGCN 0.282$\pm$.015 97.340$\pm$.143 67.312$\pm$.311 85.568$\pm$.088 73.840$\pm$.326
GatedGCN-LSPE 0.090$\pm$.001 - - - -
PNA 0.188$\pm$.004 97.940$\pm$.12 70.35$\pm$.63 - -
DGN 0.168$\pm$.003 - 72.838$\pm$.417 86.680$\pm$.034 -
GSN 0.101$\pm$.010 - - - -
CIN 0.079$\pm$.006 - - - -
CRaW1 0.085$\pm$.004 97.944$\pm$.050 69.013$\pm$.259 - -
GIN-AK+ 0.080$\pm$.001 - 72.19$\pm$.13 86.850$\pm$.057 -
SAN 0.139$\pm$.006 - - 86.581$\pm$.037 76.691$\pm$.65
Graphormer 0.122$\pm$.006 - - - -
K-Subgraph SAT 0.094$\pm$.008 - - 86.848$\pm$.037 77.856$\pm$.104
EGT 0.108$\pm$.009 98.173$\pm$.087 68.702$\pm$.409 86.821$\pm$.020 79.232$\pm$.348
Graphormer-URPE 0.086$\pm$.007 - - - -
Graphormer-GD 0.081$\pm$.009 - - - -
GPS 0.070$\pm$.004 98.051$\pm$.126 72.298$\pm$.356 86.685$\pm$.059 78.016$\pm$.180
GRIT 0.059$\pm$.002 98.108$\pm$.111 76.468$\pm$.881 87.196$\pm$.076 80.026$\pm$.277
EXPHORMER - 98.550$\pm$.039 74.69$\pm$.125 86.740$\pm$.015 78.070$\pm$.037
Eigenformer 0.077 97.142 67.316 86.680 77.262
Eigenformer (feat.) 0.089 98.362 70.194 86.738 77.456
Table 2: Test performance on a benchmark from the long-range graph benchmarks (LRGB) work [Dwivedi et al., 2022b]. Due to limited training resources, we show results for only a single training run while other baselines report mean $\pm$ standard deviation of 4 runs with different random seeds. Color codes denote top first, second, third results.
Model Peptides-func Peptides-struct
[…]
GCN 0.5930$\pm$0.0023 0.3496$\pm$0.0013
GINE 0.5498$\pm$0.0079 0.3547$\pm$0.0045
GatedGCN 0.5864$\pm$0.0035 0.3420$\pm$0.0013
GatedGCN+RWSE 0.6069$\pm$0.0035 0.3357$\pm$0.0006
Transformer+LapPE 0.6326$\pm$0.0126 0.2529$\pm$0.0016
SAN+LapPE 0.6384$\pm$0.0121 0.2683$\pm$0.0043
SAN+RWSE 0.6439$\pm$0.0075 0.2545$\pm$0.0012
GPS 0.6535$\pm$0.0041 0.2500$\pm$0.0012
GRIT 0.6988$\pm$0.0082 0.2460$\pm$0.0012
EXPHORMER 0.6527$\pm$0.0043 0.2481$\pm$0.0007
Eigenformer 0.6414 0.2599
Eigenformer (feat.) 0.6317 0.2687
(a) ZINC
(b) MNIST
(c) CIFAR10
(d) PATTERN
(e) CLUSTER
(f) Peptides-func/struct
Figure 3: (Smoothed) potential $\sigma_{k}$ vs eigenvalue $\lambda_{k}$
for adjacent and non-adjacent nodes
4 Experimental Results
We evaluate Eigenformer on five benchmarks from the widely used Benchmarking GNNs work [Dwivedi et al., 2020], viz. ZINC, MNIST, CIFAR10, PATTERN and CLUSTER. These datasets include different graph learning objectives like node classification, graph classification and graph regression. They also cover widely different domains ranging from chemistry and computer vision to synthetically generated graphs. Further, we report numbers on two benchmarks from the Long-Range Graph Benchmark (LRGB) [Dwivedi et al., 2022b] to gauge how well SAA does in capturing long-range dependencies. Further details about the experimental setup can be found in appendix B.
Baselines: We compare our method to a variety of SOTA GNN methods including (hybrid) Graph Transformers: GraphGPS Rampášek et al. [2022], GRIT Ma et al. [2023], EGT Hussain et al. [2022], SAN Kreuzer et al. [2021], Graphormer Ying et al. [2021], K-Subgraph SAT Chen et al. [2022b] and EXPHORMER Shirzad et al. [2023], MP-GNNs: GCN Kipf and Welling [2017], GAT Veličković et al. [2018], GIN Xu et al. [2019], PNA Corso et al. [2020], GatedGCN Bresson and Laurent [2018], DGN Beaini et al. [2021], GSN Bouritsas et al. [2023] and other performant models like CIN Bodnar et al. [2021], CRaW1 Tönshoff et al. [2021] and GIN-AK+ Zhao et al. [2021].
Benchmarking GNNs [Dwivedi et al., 2020]: Table 1 compares different methods under the parameter budget used in Rampášek et al. [2022] ($\sim$500K parameters for ZINC, PATTERN and CLUSTER, $\sim$100K parameters for MNIST and CIFAR10). Eigenformer routinely beats all the MP-GNNs and performs comparably to popular Graph Transformers that use different PEs like SAN, Subgraph-SAT, Graphormer, GRIT etc. Further, Eigenformer with both spectral and feature attention does marginally better for these datasets except on ZINC, where spectral attention seems to be adequate. Feature attention appears to be most beneficial for the two directed-graph datasets: MNIST and CIFAR10.
Long-range graph benchmark [Dwivedi et al., 2022b]: We also report numbers on two benchmarks from the LRGB work [Dwivedi et al., 2022b] viz. Peptides-func (multilabel graph classification) and Peptides-struct (graph regression) in Table 2. Again, Eigenformer beats all the MP-GNN models comfortably and is competitive with other Graph Transformers albeit falling short of featuring among the best performers. Interestingly, feature attention is detrimental to performance for these long-range interactions. In conjunction with the results from Table 1, our experiments suggest that feature information is not very helpful in determining how messages ought to be exchanged between nodes.
Visualization of node potentials: It is instructive to visualize the difference in the distribution of potentials (as defined in 5) between adjacent and non-adjacent nodes. Figure 3(f) shows the (smoothed, average) potential distribution w.r.t frequency for all the datasets used in our work. In general, adjacent nodes have higher potentials at the lower end of the spectrum that vary inversely with frequency. Non-adjacent nodes on the other hand, have smaller potentials which increase towards the higher end of the spectrum. We posit that these distributional differences allow the attention mechanism to learn to discern local interactions from long-range interactions.
Visualization of learned attention matrices: Figure 4 shows the adjacency matrices (1-hop: 4(a), 5-hop: 4(b)) and the shortest-path distance matrix (4(c)) for an example graph from the ZINC dataset, along with visualizations of learned attention patterns from the trained model that most closely mimic the corresponding neighborhood pattern (4(d), 4(e), 4(f)). For example, Figure 4(e) shows the attention pattern of a head from Layer 11 of the learned model, which seemingly tries to reproduce the 5-hop neighborhood interactions similar to those shown in Figure 4(b), providing empirical evidence that Eigenformer can learn k-hop interactions.
Figure 4: k-hop neighborhood, shortest-path distance and learned attention matrices for an example graph from the ZINC dataset
Limitations: A couple of open challenges remain however, including the final test performance and computational complexity of Eigenformer. With regards to performance, we believe a more extensive hyperparameter search coupled with edge-propagation might aid in closing the gap to the current SOTA in Graph Transformers. The main bottleneck w.r.t complexity is equation 7 which can be implemented using in-place operations with $O(N^{2})$ memory but in $O(N^{3})$ time. The payoff for the added complexity is increased representational capacity, as SAA can learn $k$-hop interactions for any $k\in\mathbb{Z}$ as against upper-bounded $k$-hop neighborhoods in previous works. An obvious way to reduce the time complexity to $O(N^{2}k)$ is by using a subset of the frequencies. We test this intervention on the ZINC dataset, by randomly choosing $k$ out of $N$ eigenvalues for each example in the data at each training step. Table 4 and Figure 5 show the change in final test performance with varying $k$. The test performance remains impressive even when using $\sim 40\%$ of the eigenvalues, suggesting a graceful degradation in performance with decreasing number of eigenvalues ($k$) used.
Table 3: Test MAE for the ZINC dataset for varying number (k) of eigenvalues used. N=37.
k k/N Test MAE $\downarrow$
37 1.0 0.089
30 0.81 0.097
22 0.59 0.108
15 0.41 0.133
8 0.22 0.180
Figure 5: Percentage change in MAE with decreasing number (k) of eigenvalues used
5 Related Work
Considering the importance of graph connectivity/topology information for any GNN architecture, there have been many attempts to incorporate graph inductive biases in Transformers following the introduction of Graph Transformers in Li et al. [2019].
Graph inductive biases in Transformers: Dwivedi and Bresson [2020] performs attention only over the 1-hop neighborhood of nodes in addition to the use of Laplacian eigenvectors as node PEs. With Graph-BERT, Zhang et al. [2020] proposes to train Transformers using sampled linkless (no edges) subgraphs within their local contexts, designed with an emphasis on parallelization and mainly as a pretraining technique. They use a number of PEs like the WL Absolute Role Embedding, Intimacy-based Relative Positional Embedding and Hop-based Relative Distance Embedding, in order to retain the original graph structural information in the linkless subgraphs.
Kreuzer et al. [2021] uses a Transformer encoder to create learnable PEs from a combination of eigenvalues and eigenvectors which are then fed to the main Transformer network. Dwivedi et al. [2022a] decouples the node features and the node PEs with a separate propagation step for the PEs. They consider two different initializations for the PEs: Laplacian PE (LapPE) and Random Walk PE (RWPE), and show that RWPE outperforms LapPE, indicating that learning the sign invariance ambiguity in LapPE is more difficult than not exactly having unique node representation for each node. Ying et al. [2021] proposes the use of a fixed centrality encoding to inject degree information in their Graphormer.
With GraphiT, Mialon et al. [2021] enumerates, encodes and aggregates local sub-structures such as paths of length-k using Graph convolutional kernel networks (GCKN) [Chen et al., 2020] to create node PEs. Ma et al. [2023] proposes GRIT, again using PEs initialized with random walk probabilities but makes them flexible and learnable by composing with an MLP, and provide theoretical proof of the expressivity of such a composition. Further, they inject degree information using an adaptive degree-scaler [Corso et al., 2020] in every layer, noting that attention mechanisms are invariant to node-degrees.
Structure-aware attention mechanisms: More similar to our work, there has also been research on creating structure-aware attention mechanisms in order to infuse node structural and neighborhood similarity information directly in the attention logic. Mialon et al. [2021] views self-attention as kernel smoothing and bias the attention scores using diffusion or k-step random-walk kernels. Chen et al. [2022b] argues that such a modification is deficient in filtering out structurally dissimilar nodes, and consider a more generalized kernel accounting for local substructures like k-subtrees and k-subgraphs, around each node. In a similar approach, G et al. [2023] aggregates node feature information from k-subtrees but augment their Transformer with a global node and train positional encodings using the Neighbourhood Contrastive Loss (NCE). Ying et al. [2021] proposed to inject spatial and edge information by adding learnable encodings to the attention scores before normalization, while [Menegaux et al., 2023] augments that idea further by learning a different attention score per hidden dimension in their so-called Chromatic Self-Attention (CSA). Learned edge encodings are also incorporated in Hussain et al. [2022] both as a bias to the attention score before normalization and as a sigmoid-gate after normalization.
In contrast to previous work, to the best of our knowledge, there have been no attempts to factorize the attention matrix solely into fixed spectral potentials and learnable frequency importances.
6 Conclusion
Observing the wide variety of positional encoding schemes for Graph Transformers in the literature, we asked the question whether such encodings are essential to ace graph representational learning objectives. We introduced a novel spectrum-aware attention (SAA) mechanism that incorporates structural graph inductive biases while remaining flexible enough to learn from features in the data. SAA was also shown to be invariant to sign or basis invariance issues that plagued previous PE design strategies. We proposed Eigenformer, a Graph Transformer with SAA at its core, that can be used for a variety of graph learning objectives with different task-specific heads. We empirically test the feasibility of our method through experiments on a number of common GNN benchmarks and find that it performs comparably to many popular GNN architectures employing different schemes for PE generation. We hope this work provides the impetus to the broader graph representational learning research community to pursue the development of other GNN architectures without the use of PEs.
Appendix A Proofs
A.1 Proposition 1
(1) For an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with adjacency matrix $A$ and degree matrix $D$, we have:
$L_{norm}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{norm}$
Since $L_{norm}$ is real and symmetric, it is diagonalizable and admits an eigendecomposition as follows:
$L_{norm}=U\Lambda U^{-1}$
where the columns of $U=[u_{1},u_{2},...,u_{|\mathcal{V}|}]$ are the eigenvectors of $L_{norm}$. It follows that,
$A_{norm}=I-U\Lambda U^{-1}=U(I-\Lambda)U^{-1}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}(1-\lambda_{i})$ (14)
Thus, in general we have:
$\sum_{k=0}^{m}\theta_{k}A^{k}_{norm}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}\sum_{k=0}^{m}\theta_{k}(1-\lambda_{i})^{k}$ (15)
The continuous functions $\phi_{1}(x)=x$ and $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ in 6 can hence be approximated to an arbitrary accuracy by sufficiently wide MLPs by virtue of standard universal approximation results [Hornik et al., 1989].
(2) We need to prove that there exist functions $\phi_{1}$ and $\phi_{2}$ such that $\phi_{1}\big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\big{)}\approx f(SPD^{g}[i,j])$ for every graph $g$ with adjacency matrix $A\in\mathbb{G}_{n}$ and every node-pair $(i,j)$ in $g$.
If we can ensure that we can choose $\phi_{2}$ such that it is possible to map every value $\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})$ to the unique value $f(SPD^{g}[i,j])$, then we can simply define $\phi_{1}:\mathbb{R}\rightarrow\mathbb{R}$ as the piecewise-linear continuous function:
$\phi_{1}\Big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=\phi_{1}\Big{(}\sum_{k=1}^{n}u^{g}_{k}[i]u^{g}_{k}[j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=f(SPD^{g}[i,j])$ (16)
which can be approximated by a sufficiently wide MLP. The proof then amounts to showing the existence of such a $\phi_{2}$.
Notation: With a slight abuse of notation, let $v^{g}_{ij}=[u^{g}_{1}[i]u^{g}_{1}[j],u^{g}_{2}[i]u^{g}_{2}[j],...,u^{g}_{n}[i]u^{g}_{n}[j]]^{T}$. Further denote by $\phi_{2}^{g}$ the vector $[\phi_{2}(\lambda^{g}_{1}),...,\phi_{2}(\lambda^{g}_{n})]^{T}$, so that $\sum_{k=1}^{n}u^{g}_{k}[i]u^{g}_{k}[j]\cdot\phi_{2}(\lambda^{g}_{k})={v^{g}_{ij}}^{T}\phi^{g}_{2}$.
Now, let $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ for some $m\in\mathbb{Z}$.
Case 1: $v^{g}_{ij}=v^{g}_{pq}$ for some graph $g$ and node-pairs $(i,j)$ and $(p,q)$ in $g$.
In this case, by virtue of 15, $SPD^{g}[i,j]=SPD^{g}[p,q]$ since $A^{k}_{norm}[i,j]=A^{k}_{norm}[p,q]$, which implies that there is a $k$-hop path from $i$ to $j$ iff there is a $k$-hop path from $p$ to $q$. Moreover, in this case ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$.
Case 2: $v^{g}_{ij}\neq v^{g}_{pq}$ for some graph $g$ and node-pairs $(i,j)$ and $(p,q)$ in $g$.
If $f(SPD^{g}[i,j])=f(SPD^{g}[p,q])$, we simply set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=f(SPD^{g}[i,j])$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=f(SPD^{g}[p,q])$.
If $f(SPD^{g}[i,j])\neq f(SPD^{g}[p,q])$ and ${v^{g}_{ij}}^{T}\phi^{g}_{2}\neq{v^{g}_{pq}}^{T}\phi^{g}_{2}$, we are done. If however, ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$, choose an index $r$ such that $v^{g}_{ij}[r]\neq v^{g}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g}_{ij}}^{T}\phi^{g}_{2}$ becomes unequal to ${v^{g}_{pq}}^{T}\phi^{g}_{2}$, in which case we again can set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=f(SPD^{g}[i,j])$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=f(SPD^{g}[p,q])$. We are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$ as the set of such constraints is finite and bounded.
Case 3: $v^{g_{1}}_{ij}=v^{g_{2}}_{pq}$ for some graphs $g_{1},g_{2}$ and node-pairs $(i,j)$ in $g_{1}$ and $(p,q)$ in $g_{2}$.
If $f(SPD^{g_{1}}[i,j])=f(SPD^{g_{2}}[p,q])$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$.
If $f(SPD^{g}[i,j)]\neq f(SPD^{g}[p,q])$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $\lambda^{g_{1}}_{r}\neq\lambda^{g_{2}}_{r}$ (which exists necessarily since otherwise $SPD^{g_{1}}[i,j]=SPD^{g_{2}}[p,q]$ due to 15). We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$. Similarly to Case 2 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$.
Case 4: $v^{g_{1}}_{ij}\neq v^{g_{2}}_{pq}$ for some graphs $g_{1},g_{2}$ and node-pairs $(i,j)$ in $g_{1}$ and $(p,q)$ in $g_{2}$.
If $f(SPD^{g_{1}}[i,j])=f(SPD^{g_{2}}[p,q])$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$.
If $f(SPD^{g}[i,j])\neq f(SPD^{g}[p,q])$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $v^{g_{1}}_{ij}[r]\neq v^{g_{2}}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$. Similarly to Case 2 and 3 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$.
Thus, we have proven that the continuous function $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ for some $m\in\mathbb{Z}$ can be chosen to ensure the continuity of $\phi_{1}$ defined in 16. Both functions can hence be approximated by MLPs to an arbitrary accuracy due to Hornik et al. [1989].
Appendix B Experiment Details
B.1 Description of Datasets
Details specific to each dataset used in our results are presented in Table 4.
Table 4: Overview of datasets used in our work
Dataset #graphs #nodes #edges Directed Level Task
ZINC 12,000 $\sim$23.2 $\sim$24.9 No Graph Regression
MNIST 70,000 $\sim$70.6 $\sim$564.5 Yes Graph 10-Class Classification
CIFAR10 60,000 $\sim$117.6 $\sim$941.2 Yes Graph 10-Class Classification
PATTERN 14,000 $\sim$118.9 $\sim$3039.3 No Node Binary Classification
CLUSTER 12,000 $\sim$117.2 $\sim$2150.9 No Node 6-Class Classification
Peptides-func 15,535 $\sim$150.9 $\sim$307.3 No Graph 10-Class Multilabel
Classification
Peptides-struct 15,535 $\sim$150.9 $\sim$307.3 No Graph 11-Task Regression
B.2 Hyperparameters
Due to a limited training budget we could not perform a large scale hyperparameter search and mostly borrowed configurations from Ma et al. [2023] and Rampášek et al. [2022], keeping in line with the parameter budgets, optimizer (AdamW with $betas=(0.9,0.99)$ and $eps=1e-8$) and learning rate schedule (warmup+cosine annealing) used therein. Notably, due to the simple attention mechanism of Eigenformer (without feature attention), we could use deeper/wider models for training. Further details are included in Table 5 and 6 (SSR = Signed-square-root. For the notation A/B, A=metric without feature attention and B=metric with feature attention).
Table 5: Best hyperparameter configurations for the benchmarks from Dwivedi et al. [2020]
Hyperparameter ZINC MNIST CIFAR10 PATTERN CLUSTER
# Transformer Layers 12 3 3 10 16
# Heads 8 8 8 8 8
$\psi$ SSR SSR Identity SSR SSR
Node feature embedding 128 - - - -
Edge feature embedding 128 - - - -
Hidden dim. 72/56 64/56 64/56 64/64 56/56
$\phi_{1}$ Hidden dim. 28/28 28/24 28/24 36/28 28/24
$\phi_{2}$ Hidden dim. 28/28 28/24 28/24 36/28 28/24
Dropout 0 0 0 0 0
Attention dropout 0.2 0.2 0.2 0.5 0.5
Graph pooling sum mean mean - -
Batch size 128/512 512 512 16/24 16
Learning rate 0.001 0.001 0.001 0.0005 0.0005
# Epochs 2000 200 200 100 100
# Warmup epochs 50 5 5 5 5
Weight decay 1e-5 1e-5 1e-5 1e-5 1e-5
# Parameters (w/o feat) 509849 106746 106874 476929 486006
# Parameters (w/ feat) 479481 110234 110346 472321 479734
# Epoch time (s) 21/23 198/222 793/840 571/619 550/582
# Accelerator (NVIDIA) A10G A10G A10G A10G A10G
# Peak memory (GB) 19.44/20.54 19.95/19.19 20.39/19.06 16.38/21.17 17.9/17.53
Table 6: Best hyperparameter configurations for the benchmarks from Dwivedi et al. [2022b]
Hyperparameter Peptides-func Peptides-struct
# Transformer Layers 10 10
# Heads 8 8
$\psi$ SSR SSR
Hidden dim 80/64 80/64
$\phi_{1}$ Hidden dim 32/28 32/28
$\phi_{2}$ Hidden dim 32/28 32/28
Dropout 0 0
Attention dropout 0.5 0.5
Graph pooling mean mean
Batch size 12/16 14/14
Learning rate 0.0003 0.0003
# Epochs 200 200
# Warmup epochs 5 5
Weight decay 1e-5 1e-5
# Parameters (w/o feat) 510570 510651
# Parameters (w/ feat) 510651 478779
# Epoch time (s) 1740/2736 2541/2700
# Accelerator (NVIDIA) H100 H100
# Peak memory (GB) 60.76/75.76 70.77/66.69